beamline experiments
  Beamline Logbook  Not logged in ELOG logo
Message ID: 104     Entry time: Fri Aug 6 12:40:11 2004     In reply to: 103
Author: Eric D. 
Type: Fault 
Category: General 
Subject: PSS serious faults  
Hi all, 

The P5 minihutch was opened today following a last serious fault at 8h30
this morning. They had to replace two rubber gaskets on the upstream
cylinder on the P5 actuator. Everything went well and the beam was back at
12h15. There will need to be another white beam test of the P5 during Dick's
run when we go to white beam mode. Pictures are below FYI.

Attch 1: The left cylinder is faulty. Note the top cooling rings are taken off 

Attch 2: Same as before. Note the bare Cu rod shows the side where they will
disassemble the cylinder

Attch. 3: The front cylinder limit switches. The cylinder did not have enough
pressure to make reliable contact with the limits. This picture is before repair.
The limit assembly had to be removed for replacements of the gasket. The PSS
group had to also be called in to remove the PSS system. The vacuum group (Joe
Gagliano x3098) was in charge of the seal. 

Attch.4: the replaced seal and disassembled PSS wiring and P5 cylinder.

Eric


> Hi all,
> 
> we've had PSS issues in the last day or so on 7ID. Dal-Hyun paged me today
> at 9h30 pm. He managed to get a floor coordnator to help to reset the PSS
> system. 
> It had happened to him already once before. Then around 00h30 am, while
> searching the hutch or opening the P5 shutter, a serious fault appeared on
> the PSS panel near 7ID-A. The FC reseted it, but he can only do it one more
> time tonight (serious faults cannot be overlooked too long). Anyway, someone
> will likely look at it tomorrow.
> 
> ED
Attachment 1: P0002341.JPG  336 kB
Attachment 2: P0002342.JPG  333 kB
Attachment 3: P0002343.JPG  298 kB
Attachment 4: P0002346.JPG  290 kB
ELOG V2.5.2